OpenStudy (anonymous):

A store had 250 bottles of water. Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. Which recursive function best represents the number of bottles of water in the store, given that f(0) = 250?

OpenStudy (anonymous):

(A) f(n) = f(n - 1) * 0.6 + 48, n > 0 (B) f(n) = 250 - f(n - 1) * 0.4 + 48, n > 0 (C) f(n) = f(n - 1) * 0.4 + 48, n > 0 (D) f(n) = 250 - f(n - 1) * 0.6 + 48, n > 0

OpenStudy (anonymous):

The answer is not C or B. Is it D?

OpenStudy (cwrw238):

Hint as 40% are sold this leaves 60% or 0.60

OpenStudy (anonymous):

Oh, ok. That completely rules out B and C. We are left with A and D.

OpenStudy (anonymous):

But why?

OpenStudy (cwrw238):

we start with 250 substituting n = 1,2 and 3 in B

OpenStudy (cwrw238):

try that

OpenStudy (anonymous):

ok give me some time.

OpenStudy (anonymous):

250-f(1-1)*0.4+48 = 298?

OpenStudy (cwrw238):

f(1-1) = f(0) = 250

OpenStudy (cwrw238):

so its 250 - 0.4*250 + 48 = 198

OpenStudy (anonymous):

Yeah okay. 250 - f(1-1) * 0.4 + 48 = 198

OpenStudy (cwrw238):

when n =2 what do you get

OpenStudy (anonymous):

250 - f(2-1) * 0.4 + 48 = 98? I'm not sure on this

OpenStudy (anonymous):

does f(2-1) = 100?

OpenStudy (anonymous):

or does it equal to 500? or 150?

OpenStudy (cwrw238):

no f(1) = 198

OpenStudy (anonymous):

How did u get 198?

OpenStudy (cwrw238):

we worked it out earlier one

OpenStudy (cwrw238):

actually it can't be B

OpenStudy (anonymous):

Yeah that's what I said earlier lol

OpenStudy (cwrw238):

because we dont start with 250 every week

OpenStudy (cwrw238):

yes you were right

OpenStudy (cwrw238):

what about C? dont forget f(0) = 250

OpenStudy (anonymous):

It can't be C. I put that answer before and it was incorrect.

OpenStudy (cwrw238):

the book said its incorrect?

OpenStudy (anonymous):

Not the book. the quiz on FLVS

OpenStudy (cwrw238):

yes its incorrect A is the correct answer because f(0) * 0.6 + 48 gives 150 + 48 = 198 for first week which is then input to the next week

OpenStudy (cwrw238):

I saw that earlier but got confused myself sorry

OpenStudy (anonymous):

oh it's alright lol

OpenStudy (cwrw238):

I guess I'm tired!!

OpenStudy (cwrw238):

A is definitely correct

OpenStudy (anonymous):

250 * 0.6 = 150 that's how u got that right?

OpenStudy (cwrw238):

yes - then add the 48

OpenStudy (anonymous):

Yeah add the 48. Thank you so much! Hope you feel better.

OpenStudy (cwrw238):

ty yw

OpenStudy (anonymous):

yw too